OpenStudy (anonymous):

Check my work: What is the volume of the solid obtained by rotating the region bounded by the curves y = 1/x and the x-axis between x =1 and x =2 about the x-axis?

OpenStudy (anonymous):

My work:\[\pi \int\limits_{2}^{1} (1/x)^2 -> 1/x^2 -> x ^{-2}\] \[\pi \int\limits_{2}^{1} x ^{-2} = x ^{-1}.\]

OpenStudy (anonymous):

\[\pi[1/1 - 1/2] = \pi/2\]

OpenStudy (nubeer):

well yeah looks right.

OpenStudy (phi):

close, but you lost a minus sign. I would use the limits 1 to 2 (not 2 to 1, which reverses the sign of the answer) also, the integral of x^(-2) is -x^(-1) Test: d/dx - x^-1 = +x^-2

OpenStudy (anonymous):

When I went from 1 to 2, I got -pi/2. The textbook says the answer is pi/2.

OpenStudy (phi):

That is because you did not do the integral of x^-2 correctly you should get \[ -x^{-1} \] not + x^-1

OpenStudy (anonymous):

yep, that's right! I got the right answer now! Thanks!